graph: y = x squared - 4x - 3: graph: y = x squared - 4x - 3
Answer by Alan3354(1916)

graph: y = x squared - 4x - 3
------------------
OK
Solved by pluggable solver: SOLVE quadratic equation (work shown, graph etc)
Quadratic equation ax^2+bx+c=0 (in our case 1x^2+-4x+-3 = 0) has the following solutons:

x[12] = (b+-sqrt( b^2-4ac ))/2\a

For these solutions to exist, the discriminant b^2-4ac should not be a negative number.

First, we need to compute the discriminant b^2-4ac: b^2-4ac=(-4)^2-4*1*-3=28.

Discriminant d=28 is greater than zero. That means that there are two solutions:  x[12] = (--4+-sqrt( 28 ))/2\a.

x[1] = (-(-4)+sqrt( 28 ))/2\1 = 4.64575131106459
x[2] = (-(-4)-sqrt( 28 ))/2\1 = -0.645751311064591

Quadratic expression 1x^2+-4x+-3 can be factored:
1x^2+-4x+-3 = (x-4.64575131106459)*(x--0.645751311064591)
Again, the answer is: 4.64575131106459, -0.645751311064591. Here's your graph:
graph( 500, 500, -10, 10, -20, 20, 1*x^2+-4*x+-3 )

To be more clear, the solutions are:
2 + sqrt(7) and
2 - sqrt(7)